• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name V. Sulkosky

    Log entry time 16:06:34 on February 06, 2009

    Entry number 259269


    keyword=Moved S2m Left Side ADC Channels

    After the DAQ was split into two DAQ's, the first ADC channel on the left
    side of s2m went missing in the online replay. The TDC was still there.
    I checked the signal at the patch panel upstairs and downstairs. The PMT
    signal was in the middle of the ADC gate. When I moved this signal into
    a different ADC, the second signal in the adjacent channel also vanished. I
    decided to move the entire set of 16 channels from ROC 3, Slot 25
    channels 16-31 to Slot 23, channels 32-47. These channels were previously
    used for the left side of the A1 aerogel.

    This fixed the problem, but I am concerned that the ADC in slot 25 maybe
    starting to have issues. I've also noticed some problems with channel 0
    in this ADC.

    I will update the pedestals for S2m soon.

    A copy of this log entry has been emailed to: brads,moffit,rom